A given mass of chlorine occupies 38cm cube at 20°c.Determine the final volume at 45°c pressure remains conatant
1
Expert's answer
2018-02-21T03:43:08-0500
The pressure remains constant, thus,
V_2/V_1 =T_2/T_1
V_2=V_1 T_2/T_1 =38 (45+273)/(20+273)=41 cm^3.